Group or not group? The set of Mnxn (R) of all nxn matrices under multiplication.


1
Expert's answer
2021-07-19T16:20:19-0400

Question. Group or not group? The set of Mnxn (R) of all nxn matrices under multiplication.


Answer. This set is not a group under multiplication.


Proof. Let "A\in M_{n\times n}(R)", "A=\begin{pmatrix}
 0 & 0 & \dots & 0 \\
 0 & 0 & \dots & 0 \\
 \dots & \dots & \dots & \dots \\
 0 & 0 & \dots & 0
\end{pmatrix}".

Suppose that matrix "A" is invertible (there exists an "B\in M_{n\times n}(R)", such that "A\cdot B=B\cdot A=I_n").

On the another hand, for every matrix "C\in M_{n\times n}(R)" we have "A\cdot C=C\cdot A=0".

Therefore, in the set of all matrices "n\times n" there exists a non-invertible matrix, so "M_{n\times n}(R)" isn't a group.